### v5

# Simple Policy Optimization

Abstract
PPO (Proximal Policy Optimization) algorithm has demonstrated excellent performance in many fields, and it is considered as a simple version of TRPO (Trust Region Policy Optimization) algorithm. However, the ratio clipping operation in PPO may not always effectively enforce the trust region constraints, this can be a potential factor affecting the stability of the algorithm. In this paper, we propose Simple Policy Optimization (SPO) algorithm, which introduces a novel clipping method for KL divergence between the old and current policies. Extensive experimental results in Atari 2600 environments indicate that, compared to the mainstream variants of PPO, SPO achieves better sample efficiency, extremely low KL divergence, and higher policy entropy, and is robust to the increase in network depth or complexity. More importantly, SPO maintains the simplicity of an unconstrained first-order algorithm. Our code is available at https://github.com/MyRepositories-hub/Simple-Policy-Optimization.
1 Introduction
Policy based model-free reinforcement learning algorithms have been widely applied in various fields, and have achieved performance surpassing humans including board games [1, 2, 3], video games [4, 5, 6], autonomous driving [7, 8] and intelligent control [9, 10, 11]. Although there are two main learning paradigms in deep reinforcement learning: value-based and policy-based, impressive algorithms such as PPO [12] and A3C [13] rely on the Actor-Critic (AC) framework [14], which is a synthesis of the two paradigms.
Proximal Policy Optimization (PPO) [12] algorithm implicitly limits the difference between the old policy and the current policy through the ratio clipping operation, which is reflected in the Trust Region Policy Optimization (TRPO) [15] algorithm as a hard constraint on their KL divergence expectation. Ratio clipping avoids solving the expensive KL divergence constraint problem, so the PPO algorithm can also be well extended to large-scale reinforcement learning environments [4, 6]. Prior works have discussed how ratio clipping can effectively implement trust region constraints. For example, even if the probability ratio is bounded, the corresponding KL divergence is not necessarily bounded [16]. The performance improvement of PPO algorithm is likely to mainly depend on “code-level optimizations” [17]. There are also some works that design novel probability ratio clipping methods or objective functions [18, 19, 20, 21], or alternate between a standard RL phase and a distillation phase during the training stage to improve sample efficiency of PPO [22, 23].
There are many ways to avoid excessive KL divergence, which gives birth to many variants of PPO. For example, KL divergence can be added as a penalty term to the surrogate objective and introduce adaptive penalty coefficient [12]. During the training process of PPO, a KL threshold is set, and an early stopping strategy is adopted when the KL divergence reaches the threshold [24]. Alternatively, an early stopping strategy can be implemented based on the deviation degree of probability ratio instead of KL divergence [25].
However, the above works may not have fundamentally changed the characteristics of the PPO algorithm (we acknowledge that the surrogate objective with ratio clipping remains the preferred choice in most cases), nor have they effectively constrained the KL divergence between the old and current policies. We propose Simple Policy Optimization (SPO) algorithm to compensate for these gaps, SPO does not even require the clipping of probability ratio, instead of a novel clipping method for the KL divergence between the old and current policies. Using SPO, we highlight three important observations about it:
  • It can effectively limit the average KL divergence between the current policy and the old policy in almost all environments, while the PPO algorithm inevitably results in a high value of KL divergence between the old and current policies.
  • It requires the calculation of the KL divergence between the old and current policies, which does not increase much computational overhead, and still, it maintains the simplicity of an unconstrained first-order algorithm.
  • The aforementioned characteristics of the SPO algorithm are robust to the increase of network depth or complexity. In contrast, the increase in KL divergence during the training process of the PPO algorithm becomes more evident as the network deepens.
Based on the above three points, SPO can achieve better monotonic improvement than PPO, and its first-order property makes it easily scalable to distributed training with many workers, delivering strong performance as well.
2 Background
In this section, we introduce the background knowledge of policy gradients in reinforcement learning, as well as the TRPO and PPO algorithms.
2.1 Policy Gradient
Reinforcement learning is generally defined by a tuple $(\mathcal{S},\mathcal{A},r,\mathcal{P},\rho_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state space and action space, $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the probability distribution of the state transition function, $\rho_{0}:\mathcal{S}\mapsto[0,1]$ is the initial state distribution, and $\gamma\in[0,1]$ is the discount factor.
For each trajectory
$\tau=(s_{1},a_{1},r_{1},\dots,s_{T-1},a_{T-1},r_{T-1},s_{T}),$ (1)
the objective function can be written as
$J(\theta)=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[R(\tau)\right]=\sum_{\tau}R(\tau)\cdot p_{\theta}(\tau),$ (2)
where $\theta$ is the parameters of the policy network, $R(\tau)=\sum_{i=1}^{T-1}\gamma^{i-1}r_{i}$, and the policy gradient can be estimated as
$\begin{split}\nabla J(\theta)&=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[R(\tau)\cdot\nabla\log p_{\theta}(\tau)\right]\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left\{R(\tau)\cdot\nabla\log\left[\rho_{0}(s_{1})\cdot\prod_{t=1}^{T-1}\pi_{\theta}(a_{t}|s_{t})\cdot\mathcal{P}(s_{t+1}|s_{t},a_{t})\right]\right\}\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left\{R(\tau)\cdot\nabla\left[\log\rho_{0}(s_{1})+\sum_{t=1}^{T-1}\log\pi_{\theta}(a_{t}|s_{t})+\log\mathcal{P}(s_{t+1}|s_{t},a_{t})\right]\right\}\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[\sum_{t=1}^{T-1}R(\tau)\cdot\nabla\log\pi_{\theta}(a_{t}|s_{t})\right]\approx\frac{1}{N}\sum_{n=1}^{N}\sum_{t=1}^{T_{n}-1}\left[R(\tau^{n})\cdot\nabla\log\pi_{\theta}(a_{t}^{n}|s_{t}^{n})\right],\\ \end{split}$ (3)
now we can estimate $\nabla J(\theta)$ through the Monte Carlo (MC) method [26].
2.2 Trust Region Policy Optimization
In TRPO, the following monotonic improvement theorem is given:
Theorem 1.
[15] Let $\alpha=D_{\rm TV}^{\max}(\pi,\tilde{\pi})$ and $\epsilon=\max_{s}\lvert\mathbb{E}_{a\sim\tilde{\pi}}\left[A^{\pi}(s,a)\right]\rvert$, the following bound holds:
$\eta(\tilde{\pi})\geq L_{\pi}(\tilde{\pi})-\frac{4\epsilon\gamma}{(1-\gamma)^{2}}\cdot\alpha^{2},$ (4)
where $L_{\pi}$ matches the original objective $\eta$ to first order if the policy $\pi$ is a differentiable function of a parameter vector [27].
Noting the relationship between total variation divergence and the KL divergence $D_{\rm TV}(\pi,\tilde{\pi})^{2}\leq D_{\rm KL}(\pi,\tilde{\pi})$, we have the following corollary:
Corollary 2.
Let $D_{\rm KL}^{\max}(\pi,\tilde{\pi})=\max_{s}D_{\rm KL}(\pi,\tilde{\pi})$, the following bound holds:
$\eta(\tilde{\pi})\geq L_{\pi}(\tilde{\pi})-\frac{4\epsilon\gamma}{(1-\gamma)^{2}}\cdot D_{\rm KL}^{\max}(\pi,\tilde{\pi}).$ (5)
Theoretically, this lower bound is really powerful, this is because if we denote it as
$M_{\pi}(\tilde{\pi})=L_{\pi}(\tilde{\pi})-\frac{4\epsilon\gamma}{(1-\gamma)^{2}}\cdot D_{\rm KL}^{\max}(\pi,\tilde{\pi}),$ (6)
we have $\eta(\tilde{\pi})\geq M_{\pi}(\tilde{\pi})$ and $\eta(\pi)=M_{\pi}(\pi)$, and then we obtain
$\eta(\tilde{\pi})-\eta(\pi)\geq M_{\pi}(\tilde{\pi})-M_{\pi}(\pi).$ (7)
This means that we can indirectly optimize the original objective $\eta$ by optimizing the lower bound $M_{\pi}$. Through a heuristic approximation, the maximum KL divergence $D_{\rm KL}^{\max}$ in the Corollary 2 is approximated as the average KL divergence and used as a constraint to obtain the formalization of TRPO algorithm, that is,
$\begin{split}\max_{\theta}\enspace&\mathbb{E}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left[\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t})\right]\\ \text{s.t.}\hskip 2.5pt\enspace&\mathbb{E}\left[D_{\text{KL}}(\pi_{\theta_{\rm old}},\pi_{\theta})\right]\leq\delta.\\ \end{split}$ (8)
This constrained optimization problem can be solved using the conjugate gradient algorithm followed by a line search.
2.3 Proximal Policy Optimization
PPO is an algorithm that attains the data efficiency and reliable performance of TRPO, while using only first-order optimization. PPO algorithm has two main variants, one of them is called PPO-Penalty, that is,
$\max_{\theta}\enspace\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left\{\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t})-\beta\left[D_{\text{KL}}(\pi_{\theta_{\text{old}}},\pi_{\theta})\right]\right\},$ (9)
where $\beta$ is adaptive KL penalty coefficient. Another one is called PPO-Clip, which can be formalized as
$\max_{\theta}\enspace\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\Bigg{\{}\min\Bigg{[}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t}),\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})},1\pm\epsilon\right)\cdot\hat{A}(s_{t},a_{t})\Bigg{]}\Bigg{\}}.$ (10)
PPO-Clip has been widely adopted by the academic community due to its simple form and reliable performance.
3 Method
In TRPO, trust region are imposed by limiting the expectation of the KL divergence between the old and current policies. Here, we denote the KL divergence under state $s_{t}$ as
$d=D_{\text{KL}}\left[\pi_{\theta_{\rm old}}(\cdot|s_{t}),\pi_{\theta}(\cdot|s_{t})\right]=\sum_{a\in\mathcal{A}}\pi_{\theta_{\rm old}}(a|s_{t})\cdot\log\frac{\pi_{\theta_{\rm old}}(a|s_{t})}{\pi_{\theta}(a|s_{t})},$ (11)
and the surrogate objective
$\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t}).$ (12)
According to the Corollary 2, we hope to optimize the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$ with the KL divergence $d$ as small as possible, as it optimizes the monotonic improvement lower bound in policy iterations. This minimax problem seems to be equivalent to the following form:
$\max_{\theta}\enspace\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left[\frac{\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})}{d}\right].$ (13)
However, note that if we simply using the above form as the objective for the policy network, this can lead to significant problems. Now Assume that the estimated value of the advantage $\hat{A}(s_{t},a_{t})$ is positive, the policy network would strive to minimize the KL divergence in order to maximize the objective. This tendency would result in the network preferring to make only minor adjustments to the parameters, thereby hindering its ability to learn effectively from historical data.
[…]
$\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left[\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\frac{d_{\rm clip}}{d}\right],$ (14)
where $d_{\text{clip}}=\text{clip}(d,0,d_{\max})$, $d_{\max}$ is a hyperparameter that constrains the upper bound of KL divergence. As we expected:
$\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\frac{d_{\rm clip}}{d}=\left\{\begin{aligned} &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t}),&d\leq d_{\max};\\ &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\frac{d_{\max}}{d},&d>d_{\max}.\\ \end{aligned}\right.$ (15)
[…]
$\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left[\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\left(\frac{d_{\rm clip}}{d}+\sigma-1\right)\cdot\sigma\right],$ (16)
[…]
$J(\theta)=\left\{\begin{aligned} &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\frac{d_{\text{clip}}}{d},&\enspace\hat{A}(s_{t},a_{t})>0;\\ &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\left(-\frac{d_{\text{clip}}}{d}+2\right),&\enspace\hat{A}(s_{t},a_{t})<0.\\ \end{aligned}\right.$ (17)
Furthermore, if $d\leq d_{\max}$, we have $\left(d_{\text{clip}}/d+\sigma-1\right)\cdot\sigma\equiv 1$, which means that the KL divergence satisfies the trust region constraints, and the goal $J(\theta)$ will be the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$. Otherwise,
$J(\theta)=\left\{\begin{aligned} &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\frac{d_{\max}}{d},&\enspace\hat{A}(s_{t},a_{t})>0;\\ &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\left(-\frac{d_{\max}}{d}+2\right),&\enspace\hat{A}(s_{t},a_{t})<0.\\ \end{aligned}\right.$ (18)
Figure 1: Function curves of $\left(d_{\text{clip}}/d+\sigma-1\right)\cdot\sigma$ when $d_{\max}$ is set to 0.01.
We will provide intuition about the objective (16), suppose that $d_{\max}=0.01$, and the function curves is shown in Figure 1. If $\hat{A}(s_{t},a_{t})>0$ and $d>d_{\max}$, the objective $J(\theta)$ will decrease rapidly (see curve $d_{\text{clip}}/d$), this can force the KL divergence to satisfy the trust region constraints. Otherwise, when $\hat{A}(s_{t},a_{t})<0$ and $d>d_{\max}$, the function curve $-d_{\text{clip}}/d+2$ and $d_{\text{clip}}/d$ are symmetric about the line $y=1$, and $J(\theta)$ will also decrease. Therefore, to maximize the objective $J(\theta)$, SPO will secretly optimize the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$ within the trust region $d\leq d_{\max}$ as much as possible, but it is still an unconstrained first-order algorithm.
[…]
Considering that the hyperparameter $d_{\max}$ can be set manually and even dynamically changed at any moment during the training process, SPO can benefit significantly from the Corollary 2. In summary, the characteristics mentioned above enable SPO to potentially achieve better monotonic policy improvement than PPO, and its performance depends on the value of the hyperparameter $d_{\max}$, which corresponds to the degree of conservativeness in policy iterations.
Algorithm 1 Simple Policy Optimization (SPO)
1:  Input: Policy and value network parameters $\theta,w$; threshold of KL divergence $d_{\max}$
2:  Output: Optimal policy network parameters $\theta^{*}$
3:  while not converged do
4:     # Data collection
5:     Collect data $X=\left\{(s_{i},a_{i},r_{i})\right\}_{i=1}^{N}$ using the current policy network $\pi_{\theta}$
6:     for each training epoch do
7:        # Estimate advantage function
8:        Use GAE [28] technique to estimate the advantage $\hat{A}(s_{t},a_{t})$
9:        # Calculate KL divergence
10:        For each $(s_{t},a_{t})$ calculate $d\leftarrow D_{\text{KL}}\left[\pi_{\theta_{\rm old}}(\cdot|s_{t}),\pi_{\theta}(\cdot|s_{t})\right]$ and $d_{\text{clip}}\leftarrow\text{clip}(d,0,d_{\max})$
11:        # Calculate policy loss $l_{p}$, value loss $l_{v}$ and policy entropy $l_{e}$
12:        Policy loss
$l_{p}\leftarrow\left\{\begin{aligned} &-\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t}),\enspace&\theta=\theta_{\rm old}\\ &-\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\left(\frac{d_{\rm clip}}{d}+\sigma-1\right)\cdot\sigma,\enspace&\theta\neq\theta_{\rm old}\\ \end{aligned}\right.$
and total loss $l\leftarrow l_{p}+c_{1}\cdot l_{v}-c_{2}\cdot l_{e}$
13:        # Update parameters
14:        Update parameters $\theta$ and $w$ through backpropagation
15:     end for
16:     Current policy network parameters $\theta_{\rm old}\leftarrow\theta$
17:  end while
4 Experiments
We report results on the Atari 2600 environment [29, 30]. The Atari 2600 environment is a popular benchmark in the field of reinforcement learning. It consists of a set of video game environments designed for the Atari 2600 gaming console. These environments serve as testing grounds for developing and evaluating reinforcement learning algorithms and have been widely used because of its diverse set of games, ranging from simple classics like Pong to more complex titles like SpaceInvaders and Breakout. This diversity allows researchers to assess the generalization capabilities of reinforcement learning algorithms across different tasks and complexities.
We preprocessed the Atari 2600 environment, for example, the input to the network is the stacking of the last four frames resized to $84\times 84\times 4$, reward clipping and so on [31, 32]. In all experiments, we use the hyperparameters provided in the Appendix A unless otherwise specified. For each environment in Atari 2600, we compute and visualize the standard deviation of the algorithm’s performance across 3 separate runs with different random seeds. Our experiments are divided into five parts:
  1. 1.
    Comparison to PPOs. The first part compares the performance of several PPO variants and the SPO algorithm in the Atari 2600 environment, as well as the changes in KL divergence.
  2. 2.
    Over optimization. In the second part, we increase the number of update epochs from 8 to 1000 in a certain training stage to over-optimize the respective objectives of PPO and SPO, and observe their optimization of surrogate objectives and the growth trend of KL divergence.
  3. 3.
    Increasing network complexity. In the third part, we try to increase the depth of the neural network and observe how the complexity of the network structure affects the performance of PPO and SPO in the training process.
  4. 4.
    The scalability of SPO for continuous action space. In the fourth part, we further tested the scalability of SPO in a continuous action space environment, e.g., MuJoCo [33].
  5. 5.
    Sensitivity analysis. Finally, we conducted a sensitivity analysis on the hyperparameter $d_{\max}$ of the SPO algorithm. Due to the limitation of space, please refer to the Appendix B.7 for the detailed content of this section.
Figure 2: Training curves of different PPO variants and SPO algorithm in Atari 2600 environments. Among them, PPO-C represents PPO-Clip, PPO-P represents PPO-Penalty, and PPO-E represents PPO-Clip with an early stopping strategy.
4.1 Comparison to PPOs
For the sake of experimental fairness, the policy network and value network structures of both the SPO and PPO algorithms are entirely identical, see shallow network structure in Figure 4. They share a common convolutional layer designed to compress high-dimensional image inputs into a low-dimensional representation of 512 dimensions. Subsequently, a fully connected layer is connected separately to generate outputs for the policy network (with the output dimension corresponding to the action space) and the value network (with a 1-dimensional output).
Partial experimental results are shown in Figure 2, with more detailed results provided in the Appendix B.4. Here, PPO-Penalty (PPO-P) and PPO-Clip with an early stopping strategy (PPO-E) are two main variants introduced to limit the excessive increase of KL divergence of PPO during the training process. We maintain consistency in the corresponding KL divergence thresholds for these two variants and the threshold $d_{\max}$ for SPO, all set to 0.02. It can be seen that SPO nearly achieves significantly better sample efficiency than other PPO variants in most environments, ranking second only to SPO-E in a few cases, while maintaining the lowest KL divergence throughout the entire training process (see Appendix B.4). In contrast, both PPO-C and PPO-P consistently exhibit significantly high KL divergence. However, PPO-E is also able to effectively limit KL divergence, trailing only slightly behind SPO.
We also notice that SPO exhibits the highest policy entropy during training (see Appendix B.4), which indicates that SPO has a better exploration of the environment. This reveals the fundamental difference between SPO and PPO-E, and we speculate that this may be a by-product of spontaneously limiting the KL divergence (not forcibly). In fact, we conducted a sensitivity analysis on the hyperparameter $d_{\max}$ in Appendix B.7, and we found that a lower value of $d_{\max}$ leads to a higher policy entropy.
Figure 3: We over-optimize the objectives of PPO and SPO for 1000 epochs in the early stages of training, where $d_{\max}$ is set to 0.01, 0.015, 0.02, 0.025 and 0.03, respectively.
4.2 Over Optimization
To further evaluate whether SPO can satisfy trust region constraints under extreme conditions, we select three environments and over-optimize the objectives of PPO and SPO with different $d_{\max}$ for 1000 epochs in the early stages of training, results are shown in Figure 3. Where SPO-$x$ represents the SPO algorithm with the maximum KL divergence $d_{\max}$ set to $x$, PPO-$x$ represents the PPO algorithm with the ratio clipping parameter $\epsilon$ set to $x$. We can see that even in the case of over-optimization, SPO still has a very stable KL divergence, which remains below the threshold, and the surrogate objective gradually increases as the constraints become weaker ($d_{\max}$ increases). In contrast, PPO always leads to excessively high average KL divergence and has relatively poor optimization performance for surrogate objective. Therefore, to a certain extent, this indicates that SPO optimizes the surrogate objective within the trust region as much as possible, and performs better than PPO.
Figure 4: Shallow network and deeper network structure in Atari 2600 environments.
4.3 Increasing Network Complexity
We are interested in understanding how the depth or complexity of the neural networks affects the performance of both PPO and SPO algorithms in the training process, as well as their KL divergence. This subsection aims to answer the following questions: When the network structure becomes more complex, will the ability of the SPO algorithm to limit KL divergence disappear? Will the phenomenon of PPO’s inability to limit KL divergence become more evident? The Two structures are shown in Figure 4.
We did not make any changes to the hyperparameters or other settings, except for increasing the number of network layers. PPO-1 and SPO-1 represent the use of shallow networks, while PPO-2 and SPO-2 represent the use of deeper networks. The results are shown in Figure 5. As the network becomes deeper, we can see that the KL divergence of PPO during the training process becomes larger than that in the shallow networks, sometimes leading to a sharp decline in its performance. In some environments, PPO almost loses its ability to learn effectively (e.g., Breakout and DemonAttack). In contrast, SPO demonstrates strong robustness to changes in network structure. Furthermore, we observed that the average KL divergence curves of SPO-1 and SPO-2 nearly overlap throughout the entire training process, which once again validates the effectiveness of SPO in limiting KL divergence. Additionally, this suggests that SPO is more likely to play a pivotal role in complex, large-scale neural network models, such as the Transformers [34] or DMs (Diffusion Models) [35].
Figure 5: Training curves of PPO and SPO algorithms using networks with different depths in Atari 2600 environments, where PPO-1 and SPO-1 represent the use of shallow networks, while PPO-2 and SPO-2 represent the use of deeper networks.
Figure 6: Training curves in MuJoCo environments.
4.4 The Scalability of SPO for Continuous Action Space
We further tested the scalability of the SPO algorithm in a continuous action space environment, e.g., MuJoCo [33], and surprisingly, SPO still performed exceptionally well.
Typically, in the continuous action space, we assume that the output of the policy network follows a multivariate Gaussian distribution, and we further assume that the components are mutually independent. Therefore, the KL divergence between them is the sum of the KL divergences of the individual components.
Now we assume that two random variables follow Gaussian distributions: $p(x)=\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ and $q(x)=\mathcal{N}(\mu_{2},\sigma_{2}^{2})$, then their KL divergence is
$D_{\text{KL}}(p,q)=\log\frac{\sigma_{2}}{\sigma_{1}}+\frac{(\mu_{1}-\mu_{2})^{2}+\sigma_{1}^{2}-\sigma_{2}^{2}}{2\sigma_{2}^{2}}.$ (19)
The proof see Appendix B.6. For each environment in MuJoCo, we compute and visualize the standard deviation of the algorithm’s performance across 5 separate runs with different random seeds, the results are shown in Figure 6, Table 1 and Table 2.
Table 1: Average return of the last 10% training steps in MuJoCo environments.
Algorithm Ant-v4 HalfCheetah-v4 Hopper-v4 Humanoid-v4 HumanoidStandup-v4 Walker2d-v4
SPO-0.005 $4775.95^{\downarrow 440}$ $1623.82^{\color[rgb]{1,0,0}\uparrow 16}$ $1434.74^{\downarrow 37}$ $3587.51^{\color[rgb]{1,0,0}\uparrow 2765}$ $159441.18^{\color[rgb]{1,0,0}\uparrow 8996}$ $3187.8^{\downarrow 30}$
SPO-0.01 $5066.14^{\downarrow 150}$ $2328.6^{\color[rgb]{1,0,0}\uparrow 721}$ $1377.1^{\downarrow 95}$ ${\bf 3772.78}^{\color[rgb]{1,0,0}\uparrow 2950}$ $158407.45^{\color[rgb]{1,0,0}\uparrow 7962}$ ${\bf 3508.62}^{\color[rgb]{1,0,0}\uparrow 291}$
SPO-0.03 $4628.01^{\downarrow 588}$ ${\bf 2619.4}^{\color[rgb]{1,0,0}\uparrow 1012}$ ${\bf 1632.63}^{\color[rgb]{1,0,0}\uparrow 161}$ $3168.32^{\color[rgb]{1,0,0}\uparrow 2346}$ ${\bf 159841.75}^{\color[rgb]{1,0,0}\uparrow 9396}$ $3286.5^{\color[rgb]{1,0,0}\uparrow 69}$
PPO 5216.36 1607.55 1471.81 822.57 150445.3 3217.51
Table 2: Average return throughout the entire training process in MuJoCo environments.
Algorithm Ant-v4 HalfCheetah-v4 Hopper-v4 Humanoid-v4 HumanoidStandup-v4 Walker2d-v4
SPO-0.005 $1466.71^{\downarrow 524}$ $1229.05^{\downarrow 122}$ $1046.8^{\downarrow 109}$ $1318.69^{\color[rgb]{1,0,0}\uparrow 525}$ $132000.34^{\downarrow 4209}$ $1530.99^{\downarrow 362}$
SPO-0.01 $1249.85^{\downarrow 741}$ $1705.9^{\color[rgb]{1,0,0}\uparrow 355}$ $1088.38^{\downarrow 67}$ $1457.89^{\color[rgb]{1,0,0}\uparrow 664}$ $136563.02^{\color[rgb]{1,0,0}\uparrow 354}$ $1803.99^{\downarrow 89}$
SPO-0.03 $1035.62^{\downarrow 955}$ ${\bf 1861.74}^{\color[rgb]{1,0,0}\uparrow 511}$ $1132.27^{\downarrow 23}$ ${\bf 1501.02}^{\color[rgb]{1,0,0}\uparrow 708}$ ${\bf 136769.98}^{\color[rgb]{1,0,0}\uparrow 561}$ $1788.93^{\downarrow 104}$
PPO 1991.05 1351.08 1155.63 793.51 136209.27 1893.36
From Figure 6, we can see that the sample efficiency of SPO in the Humanoid-v4 environment is much better than PPO. This may be because the environment is relatively complex, and the KL divergence of PPO becomes too large (see Appendix B.6).
As we can see from Table 1, the SPO algorithm performs better than PPO in the last 10% training steps, especially for SPO-0.03, which outperforms PPO in all environments except Ant-v4. However, according to Table 2, the average performance of the SPO algorithm decreases throughout the entire training process. Overall, this indicates that SPO is capable of achieving or even surpassing the performance of PPO in continuous control tasks (especially tasks like Humanoid-v4, which are relatively more complex), because we value the final performance of the algorithm more than the average performance of the entire training process.
5 Discussion
We propose Simple Policy Optimization (SPO), a first-order algorithm for policy optimization in reinforcement learning. Experimental results in the Atari 2600 environment demonstrate that compared to the traditional PPO algorithm, SPO significantly improves sample efficiency and effectively limits the average KL divergence between the old and current policies, while exhibiting higher policy entropy. SPO strikes a favorable balance between sample complexity, simplicity and wall-time.
In this work, we do not advocate the use of the surrogate objective with ratio clipping. We show that the ratio clipping method is not effective in limiting the average KL divergence between the old and current policies, and this phenomenon becomes more pronounced as the network deepens. In contrast, SPO achieves higher sample efficiency with lower KL divergence and exhibits robustness to the depth or complexity of the network.
We have noticed that some techniques employ reinforcement learning to allow LLMs (Large Language Models) to learn human preferences [36, 37, 38, 39], called RLHF (Reinforcement Learning from Human Feedback). A crucial step in this process is to fine-tune the LLMs using policy optimization algorithms such as PPO. To prevent the “catastrophic forgetting” of LLMs, efforts have been focused on developing new algorithms that minimize the KL divergence with the base reference model during the fine-tuning process [37, 40, 41]. It would be intriguing to apply SPO to this process, as the key difference lies in the fact that the old policy is no longer the policy network at the end of the previous training phase, but instead refers to the base reference policy.
Appendix A Hyperparameters Setting
A.1 Hyperparameters Setting in Atari 2600
Table 3: Detailed hyperparameters in Atari 2600 environments.
Algorithm PPO-Clip PPO-Penalty PPO-EarlyStop SPO
Number of actors 8 8 8 8
Horizon 128 128 128 128
Learning rate $2.5\times 10^{-4}$ $2.5\times 10^{-4}$ $2.5\times 10^{-4}$ $2.5\times 10^{-4}$
Learning rate decay Linear Linear Linear Linear
Optimizer Adam Adam Adam Adam
Total steps $1\times 10^{7}$ $1\times 10^{7}$ $1\times 10^{7}$ $1\times 10^{7}$
Batch size 1024 1024 1024 1024
Update epochs 8 8 8 8
Mini-batch size 256 256 256 256
Mini-batches 4 4 4 4
GAE parameter $\lambda$ 0.95 0.95 0.95 0.95
Discount factor $\gamma$ 0.99 0.99 0.99 0.99
[…]
Value loss coeff. $c_{1}$ 1 1 1 1
Entropy loss coeff. $c_{2}$ 0.01 0.01 0.01 0.01
[…]
A.2 Hyperparameters Setting in MuJoCo
Table 4: Detailed hyperparameters in MuJoCo environments.
Algorithm PPO SPO
Number of actors 8 8
Horizon 256 256
Learning rate $3\times 10^{-4}$ $3\times 10^{-4}$
Learning rate decay Linear Linear
Optimizer Adam Adam
Total steps $3\times 10^{6}$, $1\times 10^{7}$ $3\times 10^{6}$, $1\times 10^{7}$
Batch size 2048 2048
Update epochs 10 10
Mini-batch size 512 512
Mini-batches 4 4
GAE parameter $\lambda$ 0.95 0.95
Discount factor $\gamma$ 0.99 0.99
[…]
Value loss coeff. $c_{1}$ 0.5 0.5
Entropy loss coeff. $c_{2}$ 0.0 0.0
[…]
Appendix B More Experiment Results
B.1 Training Curves of Return
Figure 7: Training curves of return in Atari 2600 environments.
B.2 Training Curves of Average KL Divergence
Figure 8: Training curves of average KL divergence in Atari 2600 environments.
B.3 Training Curves of Policy Entropy
Figure 9: Training curves of policy entropy in Atari 2600 environments.
B.4 Comparison to PPOs
(a) Training curves of average KL divergence.
(b) Training curves of policy entropy.
Figure 10: Training curves of average KL divergence and policy entropy in Atari 2600 environments.
B.5 Increasing Network Complexity
(a) Training curves of return.
(b) Training curves of average KL divergence.
Figure 11: Training curves of return and average KL divergence in Atari 2600 environments.
Figure 12: Training curves of policy entropy in Atari 2600 environments.
B.6 The Scalability of SPO for Continuous Action Space
[…]
Proof.
The KL divergence between two given Gaussian distributions $p(x)$ and $q(x)$ is defined as
$\begin{split}&D_{\text{KL}}(p,q)\\ =&\int p(x)\cdot\log\frac{p(x)}{q(x)}\text{d}x\\ =&\int p(x)\cdot\log\left[\frac{\sigma_{2}}{\sigma_{1}}\cdot\exp\left(-\frac{(x-\mu_{1})^{2}}{2\sigma_{1}^{2}}+\frac{(x-\mu_{2})^{2}}{2\sigma_{2}^{2}}\right)\right]\text{d}x\\ =&\int p(x)\cdot\log\frac{\sigma_{2}}{\sigma_{1}}\text{d}x+\int p(x)\cdot\left(-\frac{(x-\mu_{1})^{2}}{2\sigma_{1}^{2}}+\frac{(x-\mu_{2})^{2}}{2\sigma_{2}^{2}}\right)\text{d}x\\ =&\log\frac{\sigma_{2}}{\sigma_{1}}+\frac{\sigma_{1}^{2}-\sigma_{2}^{2}}{2\sigma_{1}^{2}\sigma_{2}^{2}}\cdot\underbrace{\int x^{2}p(x)\text{d}x}_{M}+\frac{\mu_{1}\sigma_{2}^{2}-\mu_{2}\sigma_{1}^{2}}{\sigma_{1}^{2}\sigma_{2}^{2}}\cdot\underbrace{\int xp(x)\text{d}x}_{N}+\frac{\mu_{2}^{2}\sigma_{1}^{2}-\mu_{1}^{2}\sigma_{2}^{2}}{2\sigma_{1}^{2}\sigma_{2}^{2}}\cdot\underbrace{\int p(x)\text{d}x}_{=1},\\ \end{split}$ (20)
where
$M=\mathbb{E}_{x\sim p(x)}(x^{2})=\mathbb{D}_{x\sim p(x)}(x)+\mathbb{E}_{x\sim p(x)}(x)^{2}=\sigma_{1}^{2}+\mu_{1}^{2},\enspace N=\mathbb{E}_{x\sim p(x)}(x)=\mu_{1},$ (21)
so we have
$\begin{split}D_{\text{KL}}(p,q)&=\log\frac{\sigma_{2}}{\sigma_{1}}+\frac{\sigma_{1}^{2}-\sigma_{2}^{2}}{2\sigma_{1}^{2}\sigma_{2}^{2}}\cdot(\sigma_{1}^{2}+\mu_{1}^{2})+\frac{\mu_{1}\sigma_{2}^{2}-\mu_{2}\sigma_{1}^{2}}{\sigma_{1}^{2}\sigma_{2}^{2}}\cdot\mu_{1}+\frac{\mu_{2}^{2}\sigma_{1}^{2}-\mu_{1}^{2}\sigma_{2}^{2}}{2\sigma_{1}^{2}\sigma_{2}^{2}}\\ &=\log\frac{\sigma_{2}}{\sigma_{1}}+\frac{(\mu_{1}-\mu_{2})^{2}+\sigma_{1}^{2}-\sigma_{2}^{2}}{2\sigma_{2}^{2}}.\end{split}$ (22)
The trends of KL divergence and policy entropy for PPO and SPO algorithms in MuJoCo are shown in Figure 13.
(a) Training curves of average KL divergence.
(b) Training curves of policy entropy.
Figure 13: Training curves of average KL divergence and policy entropy in MuJoCo environments.
B.7 Sensitivity Analysis
We conducted a sensitivity analysis on the hyperparameter $d_{\max}$ of SPO, setting it to 0.01, 0.02, and 0.03, respectively. The experimental results are shown in Figure 14 and Figure 15.
We can see that the performance of the SPO algorithm is not highly sensitive to the selection of the hyperparameter $d_{\max}$, and the average KL divergence during the training process gradually increases as $d_{\max}$ increases. Notably, as $d_{\max}$ gradually decreases, the policy entropy of the SPO algorithm gradually increases. This result seems to be expected because the KL divergence $d$ satisfies
$d=\sum_{a\in\mathcal{A}}\pi_{\theta_{\rm old}}(a|s_{t})\cdot\log\frac{\pi_{\theta_{\rm old}}(a|s_{t})}{\pi_{\theta}(a|s_{t})}=\underbrace{H\left(\pi_{\theta_{\text{old}}}(\cdot|s_{t}),\pi_{\theta}(\cdot|s_{t})\right)}_{\text{Cross Entropy}}-\underbrace{H\left(\pi_{\theta_{\text{old}}}(\cdot|s_{t})\right)}_{\text{Policy Entropy}}.$ (23)
The cross entropy is non-negative, and it will be a good choice for the agent to reduce the KL divergence by maintaining a higher policy entropy in the training process.
We believe that there is a fundamental difference between SPO and PPO-E, the latter of which introduces an early stopping strategy to forcibly limit the KL divergence, as SPO always exhibits a higher policy entropy than PPO-E when they have a similar level of KL divergence.
(a) Error bar charts of return of SPO in Atari 2600 environments during the last 10% training steps, with their $d_{\max}$ set to 0.01, 0.02, and 0.03, respectively.
(b) Training curves of return.
Figure 14: Training curves of SPO in Atari 2600 environments, with their corresponding KL divergence thresholds $d_{\max}$ set to 0.01, 0.02, and 0.03, respectively.
(a) Training curves of average KL divergence.
(b) Training curves of policy entropy.
Figure 15: Training curves of average KL divergence and policy entropy of SPO in Atari 2600 environments, with their corresponding KL divergence thresholds $d_{\max}$ set to 0.01, 0.02, and 0.03, respectively.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and precede the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [TODO]
[…]
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [TODO]
[…]
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [TODO]
[…]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [TODO]
[…]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [TODO]
[…]
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [TODO]
[…]
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [TODO]
[…]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [TODO]
[…]
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [TODO]
[…]
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [TODO]
[…]
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [TODO]
[…]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [TODO]
[…]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [TODO]
[…]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [TODO]
[…]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [TODO]
[…]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.