### v6

# Simple Policy Optimization

Abstract
As one of the most important and influential algorithms in reinforcement learning, the Proximal Policy Optimization (PPO) algorithm has demonstrated outstanding performance across various domains. It simplifies the optimization-based importance sampling process of the Trust Region Policy Optimization (TRPO) algorithm through ratio clipping. However, this simplification with ratio clipping does not always effectively enforce trust region constraints. In this paper, we introduce an algorithm named Simple Policy Optimization (SPO), which incorporates a novel clipping method for the KL divergence between the old and new policies. Extensive experimental results in both Atari 2600 and MuJoCo environments show that, compared to PPO, SPO achieves better sample efficiency, extremely low KL divergence, and higher policy entropy, while also being robust to increases in network depth or complexity. More importantly, SPO maintains the simplicity of an unconstrained first-order algorithm. Our code is available at https://github.com/MyRepositories-hub/Simple-Policy-Optimization.
1 Introduction
Reinforcement learning has made rapid progress in recent years, becoming a cornerstone tool across a wide range of disciplines. There are two primary paradigms in deep reinforcement learning, value-based and policy-based. Policy-based model-free reinforcement learning algorithms have been extensively applied across various domains, achieving performances that surpass human capabilities in areas including board games [1, 2, 3], video games [4, 5, 6], autonomous driving [7, 8], and intelligent control systems [9, 10, 11]. Furthermore, notable algorithms like PPO [12] and A3C [13] utilize the Actor-Critic (AC) framework [14], which integrates these paradigms and serves as a foundational component in numerous fields, including games, robotics, large language models (LLMs), and more. In this paper, we focus on policy-based reinforcement learning algorithms and introduce the Simple Policy Optimization (SPO), which features a novel clipping method for importance sampling.
The Proximal Policy Optimization (PPO) algorithm [12] limits the difference between the old and new policies through a ratio clipping operation, which in the Trust Region Policy Optimization (TRPO) algorithm [15] is manifested as a hard constraint on expected KL divergence. Ratio clipping circumvents the need to solve the computationally expensive KL divergence constraint problem, thus allowing the PPO algorithm to be effectively extended to large-scale reinforcement learning environments [4, 6]. Previous studies have explored the relationship between ratio clipping and trust region constraints. For instance, even if the probability ratio is bounded, the corresponding KL divergence does not necessarily be bounded [16]. The performance improvements in the PPO algorithm are likely primarily dependent on “code-level optimizations” [17]. Additionally, some works have designed novel probability ratio clipping methods or objective functions [18, 19, 20, 21], or alternated between a standard RL phase and a distillation phase during the training stage to enhance the sample efficiency of PPO [22, 23].
There are numerous strategies to mitigate excessive KL divergence, leading to various adaptations of the PPO algorithm. For instance, KL divergence can be incorporated as a penalty term within the surrogate objective, complemented by an adaptive penalty coefficient [12]. During PPO’s training phase, a predefined KL threshold is established. An early stopping mechanism is triggered once this KL divergence threshold is reached [24]. Alternatively, the early stopping strategy can be based on the deviation in the probability ratio rather than on KL divergence itself [25].
However, the aforementioned approaches have not fundamentally altered the characteristics of the PPO algorithm (we acknowledge that the surrogate objective with ratio clipping remains the preferred choice in most cases), nor have they effectively constrained the KL divergence between the old and new policies. We propose the Simple Policy Optimization (SPO) algorithm to address these shortcomings. Notably, SPO does not require the clipping of the probability ratio; instead, it introduces a novel clipping method for the KL divergence between the old and new policies. We highlight three important contributions in SPO:
  • SPO can effectively limit the average KL divergence between the old and new policies in almost all environments, while the PPO algorithm inevitably results in a high value of KL divergence between the successive policies.
  • SPO requires the calculation of the KL divergence between successive policies, which does not increase much computational overhead, and still, it maintains the simplicity of an unconstrained first-order algorithm.
  • The aforementioned characteristics of SPO are robust to increases in network depth or complexity. In contrast, the increases in KL divergence between successive policies of PPO becomes more evident as the network deepens.
Based on the above three points, SPO can achieve better monotonic improvement than PPO, and its first-order property makes it easily scalable to distributed training with many workers, delivering strong performance as well.
2 Background
In this section, we introduce the background knowledge of policy gradients in reinforcement learning, as well as the TRPO and PPO algorithms.
2.1 Policy Gradient
Reinforcement learning is generally defined by a tuple $(\mathcal{S},\mathcal{A},r,\mathcal{P},\rho_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state space and action space, $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the probability distribution of the state transition function, $\rho_{0}:\mathcal{S}\mapsto[0,1]$ is the initial state distribution, and $\gamma\in(0,1]$ is the discount factor.
For each trajectory
$\tau=(s_{0},a_{0},r_{0},\dots,s_{T-1},a_{T-1},r_{T-1},s_{T}),$ (1)
the objective function can be written as
$J(\theta)=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[R(\tau)\right]=\sum_{\tau}R(\tau)\cdot p_{\theta}(\tau),$ (2)
where $\theta$ is the parameters of the policy network, $R(\tau)=\sum_{t=0}^{T-1}\gamma^{t}r_{t}$. Then the policy gradient can be estimated as
$\begin{split}\nabla J(\theta)&=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[R(\tau)\cdot\nabla\log p_{\theta}(\tau)\right]\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left\{R(\tau)\cdot\nabla\log\left[\rho_{0}(s_{0})\cdot\prod_{t=0}^{T-1}\pi_{\theta}(a_{t}|s_{t})\cdot\mathcal{P}(s_{t+1}|s_{t},a_{t})\right]\right\}\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left\{R(\tau)\cdot\nabla\left[\log\rho_{0}(s_{0})+\sum_{t=0}^{T-1}\log\pi_{\theta}(a_{t}|s_{t})+\log\mathcal{P}(s_{t+1}|s_{t},a_{t})\right]\right\}\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[\sum_{t=0}^{T-1}R(\tau)\cdot\nabla\log\pi_{\theta}(a_{t}|s_{t})\right]\approx\frac{1}{N}\sum_{n=1}^{N}\sum_{t=0}^{T_{n}-1}\left[R(\tau^{n})\cdot\nabla\log\pi_{\theta}(a_{t}^{n}|s_{t}^{n})\right],\\ \end{split}$ (3)
and now we can estimate $\nabla J(\theta)$ through the Monte Carlo (MC) method [26].
2.2 Trust Region Policy Optimization
In TRPO [15], the following monotonic improvement theorem is given:
Theorem 1.
Let $\alpha=D_{\rm TV}^{\max}(\pi,\tilde{\pi})$ and $\epsilon=\max_{s,a}\left|A_{\pi}(s,a)\right|$, the following bound holds:
$\eta(\tilde{\pi})\geq L_{\pi}(\tilde{\pi})-\frac{4\epsilon\gamma}{(1-\gamma)^{2}}\cdot\alpha^{2},$ (4)
where $L_{\pi}$ matches the original objective $\eta$ to first order if the policy $\pi$ is a differentiable function of a parameter vector [27].
Noting the relationship between total variation divergence and the KL divergence $D_{\rm TV}(\pi,\tilde{\pi})^{2}\leq D_{\rm KL}(\pi,\tilde{\pi})$, we have the following corollary:
Corollary 2.
Let $D_{\rm KL}^{\max}(\pi,\tilde{\pi})=\max_{s}D_{\rm KL}(\pi,\tilde{\pi})$, the following bound holds:
$\eta(\tilde{\pi})\geq L_{\pi}(\tilde{\pi})-\frac{4\epsilon\gamma}{(1-\gamma)^{2}}\cdot D_{\rm KL}^{\max}(\pi,\tilde{\pi}).$ (5)
Theoretically, this lower bound is really powerful, this is because if we denote it as
$M_{\pi}(\tilde{\pi})=L_{\pi}(\tilde{\pi})-\frac{4\epsilon\gamma}{(1-\gamma)^{2}}\cdot D_{\rm KL}^{\max}(\pi,\tilde{\pi}),$ (6)
we have $\eta(\tilde{\pi})\geq M_{\pi}(\tilde{\pi})$ and $\eta(\pi)=M_{\pi}(\pi)$, and then we obtain
$\eta(\tilde{\pi})-\eta(\pi)\geq M_{\pi}(\tilde{\pi})-M_{\pi}(\pi).$ (7)
This means that we can indirectly optimize the original objective $\eta$ by optimizing the lower bound $M_{\pi}$. Through a heuristic approximation, the maximum KL divergence $D_{\rm KL}^{\max}$ in the Corollary 2 is approximated as the average KL divergence and used as a constraint to obtain the formalization of TRPO algorithm, that is,
$\begin{split}\max_{\theta}\enspace&\mathbb{E}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left[\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t})\right],\\ \text{s.t.}\hskip 2.5pt\enspace&\mathbb{E}\left[D_{\text{KL}}(\pi_{\theta_{\rm old}},\pi_{\theta})\right]\leq\delta.\\ \end{split}$ (8)
This constrained optimization problem can be solved using the conjugate gradient algorithm followed by a line search.
2.3 Proximal Policy Optimization
PPO [12] is an algorithm that attains the data efficiency and reliable performance of TRPO, while using only first-order optimization. PPO algorithm has two main variants, one of them is called PPO-Penalty, that is,
$\max_{\theta}\enspace\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left\{\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t})-\beta\left[D_{\text{KL}}(\pi_{\theta_{\text{old}}},\pi_{\theta})\right]\right\},$ (9)
where $\beta$ is adaptive KL penalty coefficient. Another one is called PPO-Clip, which can be formalized as
$\max_{\theta}\enspace\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\Bigg{\{}\min\Bigg{[}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t}),\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})},1\pm\epsilon\right)\cdot\hat{A}(s_{t},a_{t})\Bigg{]}\Bigg{\}}.$ (10)
PPO-Clip has been widely adopted by the academic community due to its simple form and reliable performance.
3 Method
In TRPO, trust region are imposed by limiting the expectation of the KL divergence between the old and new policies. Here, we denote the KL divergence under state $s_{t}$ as
$d=D_{\text{KL}}\left[\pi_{\theta_{\rm old}}(\cdot|s_{t}),\pi_{\theta}(\cdot|s_{t})\right]=\sum_{a\in\mathcal{A}}\pi_{\theta_{\rm old}}(a|s_{t})\cdot\log\frac{\pi_{\theta_{\rm old}}(a|s_{t})}{\pi_{\theta}(a|s_{t})},$ (11)
and the surrogate objective
$\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t}).$ (12)
According to the Corollary 2, we hope to optimize the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$ with the KL divergence $d$ as small as possible, as it optimizes the monotonic improvement lower bound in policy iterations. This minimax problem seems to be equivalent to the following form:
$\max_{\theta}\enspace\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left[\frac{\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})}{d}\right].$ (13)
However, note that if we simply using the above form as the objective for the policy network, this can lead to significant problems. Now Assume that the estimated value of the advantage $\hat{A}(s_{t},a_{t})$ is positive, the policy network would strive to minimize the KL divergence in order to maximize the objective. This tendency would result in the network preferring to make only minor adjustments to the parameters, thereby hindering its ability to learn effectively from historical data.
[…]
$\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left[\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\frac{d_{\rm clip}}{d}\right],$ (14)
where $d_{\text{clip}}=\text{clip}(d,0,d_{\max})$, $d_{\max}$ is a hyperparameter that constrains the upper bound of KL divergence. As we expected:
$\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\frac{d_{\rm clip}}{d}=\left\{\begin{aligned} &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t}),&d\leq d_{\max};\\ &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\frac{d_{\max}}{d},&d>d_{\max}.\\ \end{aligned}\right.$ (15)
[…]
$\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left[\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\left(\frac{d_{\rm clip}}{d}+\sigma-1\right)\cdot\sigma\right],$ (16)
[…]
$J(\theta)=\left\{\begin{aligned} &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\frac{d_{\text{clip}}}{d},&\enspace\hat{A}(s_{t},a_{t})>0;\\ &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\left(-\frac{d_{\text{clip}}}{d}+2\right),&\enspace\hat{A}(s_{t},a_{t})<0.\\ \end{aligned}\right.$ (17)
Furthermore, if $d\leq d_{\max}$, we have $\left(d_{\text{clip}}/d+\sigma-1\right)\cdot\sigma\equiv 1$, which means that the KL divergence satisfies the trust region constraints, and the goal $J(\theta)$ will be the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$. Otherwise,
$J(\theta)=\left\{\begin{aligned} &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\frac{d_{\max}}{d},&\enspace\hat{A}(s_{t},a_{t})>0;\\ &\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\left(-\frac{d_{\max}}{d}+2\right),&\enspace\hat{A}(s_{t},a_{t})<0.\\ \end{aligned}\right.$ (18)
Figure 1: Function curves of $\left(d_{\text{clip}}/d+\sigma-1\right)\cdot\sigma$ when $d_{\max}$ is set to 0.01.
We will provide intuition about the objective (16), suppose that $d_{\max}=0.01$, and the function curves are shown in Figure 1. If $\hat{A}(s_{t},a_{t})>0$ and $d>d_{\max}$, the objective $J(\theta)$ will decrease rapidly (see curve $d_{\text{clip}}/d$), this can force the KL divergence to satisfy the trust region constraints. Otherwise, when $\hat{A}(s_{t},a_{t})<0$ and $d>d_{\max}$, the function curve $-d_{\text{clip}}/d+2$ and $d_{\text{clip}}/d$ are symmetric about the line $y=1$, and $J(\theta)$ will also decrease. Therefore, to maximize the objective $J(\theta)$, SPO will secretly optimize the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$ within the trust region $d\leq d_{\max}$ as much as possible, but it is still an unconstrained first-order algorithm.
[…]
Considering that the hyperparameter $d_{\max}$ can be set manually and even dynamically changed at any moment during the training process, SPO can benefit significantly from the Corollary 2. In summary, the characteristics mentioned above enable SPO to potentially achieve better monotonic policy improvement than PPO, and its performance depends on the value of the hyperparameter $d_{\max}$, which corresponds to the degree of conservativeness in policy iterations.
Algorithm 1 Simple Policy Optimization (SPO)
1:  Input: Policy and value network parameters $\theta,w$; threshold of KL divergence $d_{\max}$
2:  Output: Optimal policy network parameters $\theta^{*}$
3:  while not converged do
4:     # Data collection
5:     Collect data $X=\left\{(s_{i},a_{i},r_{i})\right\}_{i=1}^{N}$ using the current policy network $\pi_{\theta}$
6:     for each training epoch do
7:        # Estimate advantage function
8:        Use GAE [28] technique to estimate the advantage $\hat{A}(s_{t},a_{t})$
9:        # Calculate KL divergence
10:        For each $(s_{t},a_{t})$ calculate $d\leftarrow D_{\text{KL}}\left[\pi_{\theta_{\rm old}}(\cdot|s_{t}),\pi_{\theta}(\cdot|s_{t})\right]$ and $d_{\text{clip}}\leftarrow\text{clip}(d,0,d_{\max})$
11:        # Calculate policy loss $l_{p}$, value loss $l_{v}$ and policy entropy $l_{e}$
12:        Policy loss
$l_{p}\leftarrow\left\{\begin{aligned} &-\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t}),\enspace&\theta=\theta_{\rm old}\\ &-\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\left(\frac{d_{\rm clip}}{d}+\sigma-1\right)\cdot\sigma,\enspace&\theta\neq\theta_{\rm old}\\ \end{aligned}\right.$
and total loss $l\leftarrow l_{p}+c_{1}\cdot l_{v}-c_{2}\cdot l_{e}$
13:        # Update parameters
14:        Update parameters $\theta$ and $w$ through backpropagation
15:     end for
16:     Current policy network parameters $\theta_{\rm old}\leftarrow\theta$
17:  end while
4 Experiments
We report results on the Atari 2600 environment [29, 30]. The Atari 2600 environment is a popular benchmark in the field of reinforcement learning. It consists of a set of video game environments designed for the Atari 2600 gaming console. These environments serve as testing grounds for developing and evaluating reinforcement learning algorithms and have been widely used because of its diverse set of games, ranging from simple classics like Pong to more complex titles like SpaceInvaders and Breakout. This diversity allows researchers to assess the generalization capabilities of reinforcement learning algorithms across different tasks and complexities.
We preprocessed the Atari 2600 environment, for example, the input to the network is the stacking of the last four frames resized to $84\times 84\times 4$, reward clipping and so on [31, 32]. In all experiments, we use the hyperparameters provided in the Appendix A unless otherwise specified. For each environment in Atari 2600, we compute and visualize the standard deviation of the algorithm’s performance across three separate runs with different random seeds. Our experiments are divided into five parts:
  1. 1.
    Comparison to PPOs. The first part compares the performance of several PPO variants and the SPO algorithm in the Atari 2600 environment, as well as the changes in KL divergence.
  2. 2.
    Over optimization. In the second part, we increase the number of update epochs from 8 to 1000 in a certain training stage to over-optimize the respective objectives of PPO and SPO, and observe their optimization of surrogate objectives and the growth trend of KL divergence.
  3. 3.
    Increasing network complexity. In the third part, we try to increase the depth of the neural network and observe how the complexity of the network structure affects the performance of PPO and SPO in the training process.
  4. 4.
    The scalability of SPO for continuous action space. In the fourth part, we further tested the scalability of SPO in a continuous action space environment, e.g., MuJoCo [33].
  5. 5.
    Sensitivity analysis. Finally, we conducted a sensitivity analysis on the hyperparameter $d_{\max}$ of the SPO algorithm. Due to the limitation of space, please refer to the Appendix B.6 for the detailed content of this section.
Figure 2: Training curves of different PPO variants and SPO algorithm in Atari 2600 environments. Among them, PPO-C represents PPO-Clip, PPO-P represents PPO-Penalty, and PPO-E represents PPO-Clip with an early stopping strategy.
4.1 Comparison to PPOs
For the sake of experimental fairness, the policy network and value network structures of both the SPO and PPO algorithms are entirely identical, see shallow network structure in Figure 4. They share a common convolutional layer designed to compress high-dimensional image inputs into a low-dimensional representation of 512 dimensions. Subsequently, a fully connected layer is connected separately to generate outputs for the policy network (with the output dimension corresponding to the action space) and the value network (with a 1-dimensional output).
Partial experimental results are shown in Figure 2, with more detailed results provided in the Appendix B.4. Here, PPO-Penalty (PPO-P) and PPO-Clip with an early stopping strategy (PPO-E) are two main variants introduced to limit the excessive increase of KL divergence of PPO during the training process. We maintain consistency in the corresponding KL divergence thresholds for these two variants and the threshold $d_{\max}$ for SPO, all set to 0.02. It can be seen that SPO nearly achieves significantly better sample efficiency than other PPO variants in most environments, ranking second only to SPO-E in a few cases, while maintaining the lowest KL divergence throughout the entire training process (see Appendix B.4). In contrast, both PPO-C and PPO-P consistently exhibit significantly high KL divergence. However, PPO-E is also able to effectively limit KL divergence, trailing only slightly behind SPO.
We also notice that SPO exhibits the highest policy entropy during training (see Appendix B.4), which indicates that SPO has a better exploration of the environment. This reveals the fundamental difference between SPO and PPO-E, and we speculate that this may be a by-product of spontaneously limiting the KL divergence (not forcibly). In fact, we conducted a sensitivity analysis on the hyperparameter $d_{\max}$ in Appendix B.6, and we found that a lower value of $d_{\max}$ leads to a higher policy entropy.
Figure 3: We over-optimize the objectives of PPO and SPO for 1000 epochs in the early stages of training, where $d_{\max}$ is set to 0.01, 0.015, 0.02, 0.025 and 0.03, respectively.
4.2 Over Optimization
To further evaluate whether SPO can satisfy trust region constraints under extreme conditions, we select three environments and over-optimize the objectives of PPO and SPO with different $d_{\max}$ for 1000 epochs in the early stages of training, results are shown in Figure 3. Where SPO-$x$ represents the SPO algorithm with the maximum KL divergence $d_{\max}$ set to $x$, PPO-$x$ represents the PPO algorithm with the ratio clipping parameter $\epsilon$ set to $x$. We can see that even in the case of over-optimization, SPO still has a very stable KL divergence, which remains below the threshold, and the surrogate objective gradually increases as the constraints become weaker ($d_{\max}$ increases). In contrast, PPO always leads to excessively high average KL divergence and has relatively poor optimization performance for surrogate objective. Therefore, to a certain extent, this indicates that SPO optimizes the surrogate objective within the trust region as much as possible, and performs better than PPO.
Figure 4: Shallow network and deeper network structures in Atari 2600 environments.
4.3 Increasing Network Complexity
We are interested in understanding how the depth or complexity of the neural networks affects the performance of both PPO and SPO algorithms in the training process, as well as their KL divergence. This subsection aims to answer the following questions: When the network structure becomes more complex, will the ability of the SPO algorithm to limit KL divergence disappear? Will the phenomenon of PPO’s inability to limit KL divergence become more evident? The Two structures are shown in Figure 4.
We did not make any changes to the hyperparameters or other settings, except for increasing the number of network layers. PPO-1 and SPO-1 represent the use of shallow networks, while PPO-2 and SPO-2 represent the use of deeper networks. The results are shown in Figure 5. As the network becomes deeper, we can see that the KL divergence of PPO during the training process becomes larger than that in the shallow networks, sometimes leading to a sharp decline in its performance. In some environments, PPO almost loses its ability to learn effectively (e.g., Breakout and DemonAttack). In contrast, SPO demonstrates strong robustness to changes in network structure. Furthermore, we observed that the average KL divergence curves of SPO-1 and SPO-2 nearly overlap throughout the entire training process, which once again validates the effectiveness of SPO in limiting KL divergence. Additionally, this suggests that SPO is more likely to play a pivotal role in complex, large-scale neural network models, such as the Transformer [34] or DMs (Diffusion Models) [35].
Figure 5: Training curves of PPO and SPO algorithms using networks with different depths in Atari 2600 environments, where PPO-1 and SPO-1 represent the use of shallow networks, while PPO-2 and SPO-2 represent the use of deeper networks.
Figure 6: Training curves in MuJoCo with deeper networks.
4.4 The Scalability of SPO for Continuous Action Space
We further tested the scalability of the SPO algorithm in a continuous action space environment, e.g., MuJoCo [33], and surprisingly, SPO still performed exceptionally well.
Typically, in the continuous action space, we assume that the output of the policy network follows a multivariate Gaussian distribution, and we further assume that the components are mutually independent. Therefore, the KL divergence between them is the sum of the KL divergences of the individual components.
Now we assume that two random variables follow Gaussian distributions: $p(x)=\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ and $q(x)=\mathcal{N}(\mu_{2},\sigma_{2}^{2})$, then their KL divergence is
$D_{\text{KL}}(p,q)=\log\frac{\sigma_{2}}{\sigma_{1}}+\frac{(\mu_{1}-\mu_{2})^{2}+\sigma_{1}^{2}-\sigma_{2}^{2}}{2\sigma_{2}^{2}}.$ (19)
The details see Appendix B.7. For each environment in MuJoCo, we compute and visualize the standard deviation of the algorithm’s performance across five separate runs with different random seeds, and we also employ two different network architectures with varying depths (see Figure 16). The results are shown in Figure 6, Table 1, Table 2 and Figure 7.
Table 1: Average return of the last 10% training steps with shallow networks in MuJoCo.
Algorithm Ant-v4 HalfCheetah-v4 Hopper-v4 Humanoid-v4 HumanoidStandup-v4 Walker2d-v4
PPO 5216.4 1607.5 1471.8 822.6 150445.3 3217.5
SPO 4875.4 2459.9 1402.6 3228.6 163008.1 3520.9
Table 2: Average return of the last 10% training steps with deeper networks in MuJoCo.
Algorithm Ant-v4 HalfCheetah-v4 Hopper-v4 Humanoid-v4 HumanoidStandup-v4 Walker2d-v4
PPO 527.4 2781.0 1182.6 576.7 136717.4 2512.4
SPO 2679.8 3276.0 1390.8 3231.8 152980.9 3800.0
Figure 7: Error bar chart (from left to right): PPO, SPO-0.01, SPO-0.02, and SPO-0.03 with deeper networks.
Figure 6 and Table 2 clearly show that SPO outperforms PPO in sample efficiency across all environments when training with deeper neural networks. Furthermore, as can be seen from Figure 7, the KL divergence of SPO remains controlled during the training process, whereas the KL divergence of PPO becomes significantly large, particularly in Ant-v4 and Humanoid-v4 (the gray line in Figure 7 corresponds to a value of 0.02). This divergence led to the collapse of PPO’s performance in these environments.
Due to the limitation of space, please refer to Appendix B.7 for more detailed experimental results. In summary, we speculate that the performance of the PPO algorithm mainly depends on two factors: “code-level optimizations” and a relatively simple network structure. This is why, in commonly used benchmarks, the policy and value networks of PPO are always designed to be relatively simple. While this avoids excessive KL divergence, it also limits the algorithm’s capacity to encode complex tasks.
5 Discussion
We propose Simple Policy Optimization (SPO), an unconstrained first-order algorithm for policy optimization. Experiments on the both Atari 2600 and MuJoCo suite demonstrate the utility of the approach, as SPO is better at constraining the KL divergence between successive policies, while also being robust to increases in network depth or complexity. SPO strikes a favorable balance between sample complexity, simplicity and wall-time.
In this work, we do not advocate the use of the surrogate objective with ratio clipping. We show that the ratio clipping method is not effective in limiting the KL divergence between the successive policies, this phenomenon becomes more pronounced as the network deepens. In contrast, SPO achieves higher sample efficiency with lower KL divergence and exhibits robustness to the depth or complexity of the network.
Recently, some techniques employ reinforcement learning to allow LLMs (Large Language Models) to learn human preferences [36, 37, 38, 39], called RLHF (Reinforcement Learning from Human Feedback). A crucial step in this process is to fine-tune the LLMs using policy optimization algorithms such as PPO. To prevent the “catastrophic forgetting” of LLMs, efforts have been focused on developing new algorithms that minimize the KL divergence with the base reference model during the fine-tuning process [37, 40, 41]. Given the high parameter complexity of LLMs and the shortcomings of PPO in constraining KL divergence, it will be interesting to explore the application of SPO in RLHF in the future.
Appendix A Hyperparameters Setting
A.1 Hyperparameters Setting in Atari 2600
Table 3: Detailed hyperparameters in Atari 2600 environments.
Algorithm PPO-Clip PPO-Penalty PPO-EarlyStop SPO
Number of actors 8 8 8 8
Horizon 128 128 128 128
Learning rate $2.5\times 10^{-4}$ $2.5\times 10^{-4}$ $2.5\times 10^{-4}$ $2.5\times 10^{-4}$
Learning rate decay Linear Linear Linear Linear
Optimizer Adam Adam Adam Adam
Total steps $1\times 10^{7}$ $1\times 10^{7}$ $1\times 10^{7}$ $1\times 10^{7}$
Batch size 1024 1024 1024 1024
Update epochs 8 8 8 8
Mini-batch size 256 256 256 256
Mini-batches 4 4 4 4
GAE parameter $\lambda$ 0.95 0.95 0.95 0.95
Discount factor $\gamma$ 0.99 0.99 0.99 0.99
[…]
Value loss coeff. $c_{1}$ 1 1 1 1
Entropy loss coeff. $c_{2}$ 0.01 0.01 0.01 0.01
[…]
A.2 Hyperparameters Setting in MuJoCo
Table 4: Detailed hyperparameters in MuJoCo environments.
Algorithm PPO SPO
Number of actors 8 8
Horizon 256 256
Learning rate $3\times 10^{-4}$ $3\times 10^{-4}$
Learning rate decay Linear Linear
Optimizer Adam Adam
Total steps $3\times 10^{6}$, $1\times 10^{7}$ $3\times 10^{6}$, $1\times 10^{7}$
Batch size 2048 2048
Update epochs 10 10
Mini-batch size 512 512
Mini-batches 4 4
GAE parameter $\lambda$ 0.95 0.95
Discount factor $\gamma$ 0.99 0.99
[…]
Value loss coeff. $c_{1}$ 0.5 0.5
Entropy loss coeff. $c_{2}$ 0.0 0.0
[…]
Appendix B More Experiment Results
B.1 Training Curves of Return
Figure 8: Training curves of return in Atari 2600 environments.
B.2 Training Curves of Average KL Divergence
Figure 9: Training curves of average KL divergence in Atari 2600 environments.
B.3 Training Curves of Policy Entropy
Figure 10: Training curves of policy entropy in Atari 2600 environments.
B.4 Comparison to PPOs
(a) Training curves of average KL divergence.
(b) Training curves of policy entropy.
Figure 11: Training curves of average KL divergence and policy entropy in Atari 2600 environments.
B.5 Increasing Network Complexity
(a) Training curves of return.
(b) Training curves of average KL divergence.
Figure 12: Training curves of return and average KL divergence in Atari 2600 environments.
Figure 13: Training curves of policy entropy in Atari 2600 environments.
B.6 Sensitivity Analysis
We conducted a sensitivity analysis on the hyperparameter $d_{\max}$ of SPO, setting it to 0.01, 0.02, and 0.03, respectively. The experimental results are shown in Figure 14 and Figure 15.
We can see that the performance of the SPO algorithm is not highly sensitive to the selection of the hyperparameter $d_{\max}$, and the average KL divergence during the training process gradually increases as $d_{\max}$ increases. Notably, as $d_{\max}$ gradually decreases, the policy entropy of the SPO algorithm gradually increases. This result seems to be expected because the KL divergence $d$ satisfies
$d=\sum_{a\in\mathcal{A}}\pi_{\theta_{\rm old}}(a|s_{t})\cdot\log\frac{\pi_{\theta_{\rm old}}(a|s_{t})}{\pi_{\theta}(a|s_{t})}=\underbrace{H\left(\pi_{\theta_{\text{old}}}(\cdot|s_{t}),\pi_{\theta}(\cdot|s_{t})\right)}_{\text{Cross Entropy}}-\underbrace{H\left(\pi_{\theta_{\text{old}}}(\cdot|s_{t})\right)}_{\text{Policy Entropy}}.$ (20)
The cross entropy is non-negative, and it will be a good choice for the agent to reduce the KL divergence by maintaining a higher policy entropy in the training process.
We believe that there is a fundamental difference between SPO and PPO-E, the latter of which introduces an early stopping strategy to forcibly limit the KL divergence, as SPO always exhibits a higher policy entropy than PPO-E when they have a similar level of KL divergence.
(a) Error bar charts of return of SPO in Atari 2600 environments during the last 10% training steps, with their $d_{\max}$ set to 0.01, 0.02, and 0.03, respectively.
(b) Training curves of return.
Figure 14: Training curves of SPO in Atari 2600 environments, with their corresponding KL divergence thresholds $d_{\max}$ set to 0.01, 0.02, and 0.03, respectively.
From Figure 14 we can see that the performance of the SPO algorithm is not highly sensitive to the selection of the hyperparameter $d_{\max}$.
(a) Training curves of average KL divergence.
(b) Training curves of policy entropy.
Figure 15: Training curves of average KL divergence and policy entropy of SPO in Atari 2600 environments, with their corresponding KL divergence thresholds $d_{\max}$ set to 0.01, 0.02, and 0.03, respectively.
From Figure 15 we can see that the average KL divergence during the training process gradually increases as $d_{\max}$ increases. Notably, as $d_{\max}$ gradually decreases, the policy entropy of the SPO algorithm gradually increases. However, we surprisingly found that the change trend of policy entropy in MuJoCo (see Appendix B.7) was opposite to that in Atari 2600 with the change of $d_{\max}$.
B.7 The Scalability of SPO for Continuous Action Space
[…]
Proof.
The KL divergence between two given Gaussian distributions $p(x)$ and $q(x)$ is defined as
$\begin{split}&D_{\text{KL}}(p,q)\\ =&\int p(x)\cdot\log\frac{p(x)}{q(x)}\text{d}x\\ =&\int p(x)\cdot\log\left[\frac{\sigma_{2}}{\sigma_{1}}\cdot\exp\left(-\frac{(x-\mu_{1})^{2}}{2\sigma_{1}^{2}}+\frac{(x-\mu_{2})^{2}}{2\sigma_{2}^{2}}\right)\right]\text{d}x\\ =&\int p(x)\cdot\log\frac{\sigma_{2}}{\sigma_{1}}\text{d}x+\int p(x)\cdot\left(-\frac{(x-\mu_{1})^{2}}{2\sigma_{1}^{2}}+\frac{(x-\mu_{2})^{2}}{2\sigma_{2}^{2}}\right)\text{d}x\\ =&\log\frac{\sigma_{2}}{\sigma_{1}}+\frac{\sigma_{1}^{2}-\sigma_{2}^{2}}{2\sigma_{1}^{2}\sigma_{2}^{2}}\cdot\underbrace{\int x^{2}p(x)\text{d}x}_{M}+\frac{\mu_{1}\sigma_{2}^{2}-\mu_{2}\sigma_{1}^{2}}{\sigma_{1}^{2}\sigma_{2}^{2}}\cdot\underbrace{\int xp(x)\text{d}x}_{N}+\frac{\mu_{2}^{2}\sigma_{1}^{2}-\mu_{1}^{2}\sigma_{2}^{2}}{2\sigma_{1}^{2}\sigma_{2}^{2}}\cdot\underbrace{\int p(x)\text{d}x}_{1},\\ \end{split}$ (21)
where
$M=\mathbb{E}_{x\sim p(x)}(x^{2})=\mathbb{D}_{x\sim p(x)}(x)+\mathbb{E}_{x\sim p(x)}(x)^{2}=\sigma_{1}^{2}+\mu_{1}^{2},\enspace N=\mathbb{E}_{x\sim p(x)}(x)=\mu_{1},$ (22)
so we have
$\begin{split}D_{\text{KL}}(p,q)&=\log\frac{\sigma_{2}}{\sigma_{1}}+\frac{\sigma_{1}^{2}-\sigma_{2}^{2}}{2\sigma_{1}^{2}\sigma_{2}^{2}}\cdot(\sigma_{1}^{2}+\mu_{1}^{2})+\frac{\mu_{1}\sigma_{2}^{2}-\mu_{2}\sigma_{1}^{2}}{\sigma_{1}^{2}\sigma_{2}^{2}}\cdot\mu_{1}+\frac{\mu_{2}^{2}\sigma_{1}^{2}-\mu_{1}^{2}\sigma_{2}^{2}}{2\sigma_{1}^{2}\sigma_{2}^{2}}\\ &=\log\frac{\sigma_{2}}{\sigma_{1}}+\frac{(\mu_{1}-\mu_{2})^{2}+\sigma_{1}^{2}-\sigma_{2}^{2}}{2\sigma_{2}^{2}}.\end{split}$ (23)
The shallow network and deeper network structures used in MuJoCo are shown in Figure 16.
Figure 16: Shallow network (left) and deeper network (right) structures in MuJoCo environments.
Table 5: Average return of the last 10% training steps with shallow networks in MuJoCo.
Algorithm Ant-v4 HalfCheetah-v4 Hopper-v4 Humanoid-v4 HumanoidStandup-v4 Walker2d-v4
SPO-0.01 5066.1 2328.6 1377.1 3772.8 158407.4 3508.6
SPO-0.02 4875.4 2459.9 1402.6 3228.6 163008.1 3520.9
SPO-0.03 4628.0 2619.4 1632.6 3168.3 159841.8 3286.5
PPO 5216.4 1607.5 1471.8 822.6 150445.3 3217.5
Table 6: Average return of the last 10% training steps with deeper networks in MuJoCo.
Algorithm Ant-v4 HalfCheetah-v4 Hopper-v4 Humanoid-v4 HumanoidStandup-v4 Walker2d-v4
SPO-0.01 3376.3 2336.2 1532.3 3499.8 151760.9 3016.9
SPO-0.02 2679.8 3276.0 1390.8 3231.8 152980.9 3800.0
SPO-0.03 1504.2 2855.7 1379.8 3285.8 147222.7 3260.6
PPO 527.4 2781.0 1182.6 576.7 136717.4 2512.4
(a) Error bar chart of return with shallow networks.
(b) Error bar chart of return with deeper networks.
Figure 17: Error bar chart (from left to right): PPO, SPO-0.01, SPO-0.02, and SPO-0.03.
(a) Ant-v4.
(b) HalfCheetah-v4.
(c) Hopper-v4.
(d) Humanoid-v4.
(e) HumanoidStandup-v4.
(f) Walker2d-v4.
Figure 18: We plotted the trend of how the return varied with $d_{\max}$ during the last 10% training steps.
We recorded and plotted the mean and standard deviation of the returns in the last 10% steps of the training process when SPO and PPO algorithms were respectively applied to shallow networks and deeper networks. The results are shown in Table 5, Table 6, Figure 17 and Figure 18.
(a) Training curves of return.
(b) Training curves of average KL divergence.
(c) Training curves of policy entropy.
Figure 19: Training curves of return, average KL divergence and policy entropy in MuJoCo with shallow networks.
The training curves for PPO and SPO algorithms in MuJoCo environments with shallow networks are shown in Figure 19.
(a) Training curves of return.
(b) Training curves of average KL divergence.
(c) Training curves of policy entropy.
Figure 20: Training curves of return, average KL divergence and policy entropy in MuJoCo with deeper networks.
The training curves for PPO and SPO algorithms in MuJoCo environments with deeper networks are shown in Figure 20.